When a uniform solid sphere of mass m rolls from rest down an inclined plane such that it's centre of mass descends through a height h it's translational kinetic energy will be 
1.)mgh.       2.)5/7mgh. 3.)2/7mgh 4.)mgh/2 
Hint:the correct option is 2.)2/7mgh 

Grade:11

1 Answers

Vikas TU
4 years ago
Dear stuent 
This is pure rolling 
KR/Kr = 2/5 (solid sphere)
KR = 2/7 (Total KE)
= 2/7 (mgh)
Good Luck